1.Given a tax rate of $0.0824 and an assessed valuation of $74,900, the total property tax due is:

a.

$6,171.76

b.

None of these

c.

$6,111.67

d.

$6,071.67

e.

$6,110.67

2.Given a tax rate of $0.8231 and total property tax due of $12,510, the total assessed valuation rounded to the nearest dollar is:

a.

$1,519

b.

$5,199

c.

None of these

d.

$15,199

e.

$150,199

3.

  1. Lee's toy store is worth $400,000 and is insured for $300,000. Assuming an 80% coinsurance clause and a fire that caused $190,000 of damage, the liability of the insurance company is:

    a.

    $124,500

    b.

    $142,500

    c.

    $178,125

    d.

    $187,125

    e.

    None of these
Answer #1

Property tax = Assessed Value*Tax Rate

= 74,900*0.0824

= $6,171.76

i.e. a

2.Assessed Valuation = 12,510/0.8231

= $15,199

i.e. d.

3.Liability of Insurance company = 190,000*300,000/400,000

= $142,500

i.e. b